# Logical Consistency of Large Language Models in Fact-Checking

### v1

@@ Abstract
In recent years, large language models (LLMs) have demonstrated significant success in performing varied natural language tasks such as language translation, question-answering, summarizing, fact-checking, etc. Despite LLMs’ impressive ability to generate human-like texts, LLMs are infamous for their inconsistent responses – a meaning-preserving change in the input query results in an inconsistent response and attributes to vulnerabilities of LLMs such as hallucination, jailbreaking, etc. Consequently, existing research focuses on simple para-phrasing-based consistency assessment of LLMs, and ignores complex queries that necessitates an even better understanding of logical reasoning by an LLM. Our work therefore addresses the logical inconsistency of LLMs under complex logical queries with primitive logical operators, e.g., negation, conjunction, and disjunction. As a test bed, we consider retrieval-augmented LLMs on a fact-checking task involving propositional logic queries from real-world knowledge graphs (KGs). Our contributions are three-fold. Benchmark: We introduce three logical fact-checking datasets over KGs for community development towards logically consistent LLMs. Assessment: We propose consistency measures of LLMs on propositional logic queries as input and demonstrate that existing LLMs lack logical consistency, specially on complex queries. Improvement: We employ supervised fine-tuning to improve the logical consistency of LLMs on the complex fact-checking task with KG contexts.
@@ 1 Introduction
Large language models (LLMs), e.g., ChatGPT (OpenAI, 2022), PaLM (Chowdhery et al., 2023), LLaMA (Touvron et al., 2023) permeate natural language processing with their remarkable capacity to comprehend and generate human-like texts (Min et al., 2023; Kamalloo et al., 2023). As such, there is a host of applications of LLMs in high-stake domains such as healthcare (Singhal et al., 2023; Wang et al., 2023b; Dash et al., 2023), finance (Wu et al., 2023; Yang et al., 2023a), law (Xiao et al., 2021), and education (Kasneci et al., 2023). For instance, by implicitly storing pre-trained/fine-tuned knowledge, LLMs are increasingly applied in fact-checking and question-answering tasks (Saeed et al., 2023; Quelle & Bovet, 2024).
@@
Consistency of LLM Response. Despite the success story, a notable challenge in LLMs is the inconsistency of generated responses. Consistency of an LLM – the invariance of its output under meaning-preserving alternations in its input – is a highly desirable property (Elazar et al., 2021; Liu et al., 2023c). Existing works mostly focus on semantically similar sentences (paraphrases) for consistency assessment. For example,  Kuhn et al. (2023) measure the consistency of LLMs by computing semantic uncertainty of output responses under paraphrased inputs.  Elazar et al. (2021) consider quasi-paraphrases in a cloze-style input format and assesses masked language models (Devlin et al., 2018) for consistency (details in the Appendix B).
@@
Thomas Mann $\mid$ award winner $\mid$ Nobel Prizein LiteratureFactConsider the context: \$Context.Is the fact correct given the context? \$Fact.LLMQuery (Prompt)Nobel Prizein LiteratureT.S.EliotawardwinnerAldousHuxleyawardnomineeHenriBergsonawardwinnerThomasMannawardwinnerFranzKafkainfluencedLeoTolstoyinfluenced byKnowledge graphT.S. Eliot $\mid$ award winner $\mid$ Nobel Prize in Literature Thomas Mann $\mid$ influenced $\mid$ Franz Kafka Thomas Mann $\mid$ award winner $\mid$ Nobel prize in Literature ContextBFS startingfrom Thomas Mann
Figure 1: Our LLM-based fact-checking framework with the KG context. The LLMQuery shown is an example of zero-shot prompt, that is, it does not contain examples or demonstrations.
@@
Existing works focus on the consistency of an LLM on simple queries and do not assess consistency on complex queries requiring improved logical reasoning by LLMs. Therefore, an unexplored research question is to systematically assess and improve the logical consistency of LLMs in complex logic queries with primitive logical operators: negation ($\neg$), conjunction ($\wedge$), and disjunction ($\vee$). Informally, an LLM is logically consistent to a query under negation operator if the LLM response to the base query and the negated query are semantically opposite and complies to the semantic of negation. Subsequently, an LLM is consistent on a conjunction query with (at least) two sub-queries if the individual responses to each sub-queries are logically consistent with the conjunction query. Building on logical semantics, we naturally extend the consistency assessment of LLMs to propositional logic queries and logic rules like commutative, distributive, and associative rules (Boole, 1847). Our hypothesis is that logically consistent LLMs are essential in trustworthy systems (Sun et al., 2024b): When an LLM is consistent, one can directly verify whether the LLM is right or wrong without the need to have a sophisticated benchmark, which provides the end users confidence in applying the LLM in their own workload.
@@
Benchmark. We address the lack of a logical fact-checking dataset (LFC) over KGs by introducing three datasets: FreebaseLFC, NELLLFC, and WikiLFC5). They are derived from their KG counterparts, Freebase (Bordes et al., 2013), NELL (Carlson et al., 2010), and WikiKG90Mv2 (Hu et al., 2021b), respectively. Freebase contains real-world entities such as people, places, institutions, and their relations from the Freebase KG. NELL is a dataset built from the Web via an intelligent agent called Never-Ending Language Learner. NELL has accumulated over 100 million candidate triplets by reading the web (April 2018) with different levels of confidence. WikiKG90Mv2 is a KG extracted from the entire Wikidata knowledge base.
@@
Assessment of Logical Consistency. We propose logical consistency measures of LLMs on propositional logic-based queries, which cover fact-checking queries over KGs using retrieval-augmented LLMs (§3). The queried facts involve one or more triplets connected by logical operators such as negation ($\neg$), conjunction ($\wedge$), and disjunction ($\vee$). For consistency assessment, we consider a variety of scenarios, such as both simple and complex facts, diverse logic rules, and LLMQuery with and without KG contexts (§5). We demonstrate that existing LLMs such as Llama2222-$7$B, Llama2222-$13$B, and Gemma-$2$B, while being more accurate with KG contexts, are not consistent when considering logical equivalents or variants of these facts.
Improvement of Logical Consistency. We demonstrate how existing LLMs can be fine-tuned to improve their consistency via supervised fine-tuning (Radford et al., 2019; Devlin et al., 2018)4). We showcase that pre-train, prompt, and predict paradigm is often insufficient to improve the consistency of LLMs for complex fact-checking with KG contexts, thus we resort to pre-train, fine-tune, and predict approach. Although supervised fine-tuning is shown to be effective in other downstream tasks (Wei et al., 2021; Li & Liang, 2021), it has not yet been conducted for logical fact-checking queries over KGs. Furthermore, fine-tuning is feasible in our case due to the availability of abundant training data with ground-truth labels from the KG (§4.2). Additionally, we leverage parameter-efficient fine-tuning based on QLoRA (Quantized Low-Rank Adaptation) (Dettmers et al., 2024), thus improving the fine-tuning efficiency.
Experimental results show that the fine-tuned LLMs improve the consistency by 14% on average, while our optimization techniques scale both fine-tuning and inferencing to large KGs and complex fact-checking tasks. Our experiments also show that models fine-tuned on simple facts generalize well to more complex, unseen facts and rules (§5). This is possible due to complex facts and rules being decomposable into a collection of simple facts (Proposition 1). As ablation studies, we analyze the impact of different KG retrieval methods to build the KG context in LLMQuery, and the impact of context length and different modes of prompting such as zero-shot, few-shot, and chain-of-thought on assessing and improving the logical consistency of LLMs.
@@ 2 Background
We provide preliminaries on knowledge graphs and propositional logic, which constitute the background for assessing logical consistency in LLMs. Additional related work is in the Appendix B
@@ 3 Measuring Logical Consistency on Propositional Logic Facts
In this section, we motivate and define the quantitative measure of logical consistency of an LLM response on propositional logic facts, presented as LLMQueries (Figure 1). When prompted the LLM with a fact $q$, the LLM either accepts $q$ as a valid fact or rejects it, that is, $\mathtt{LLM}(q)\in\{0,1\}$, which is a binary function that either validates $q$ with response $1$, or invalidates $q$ with response 00. In this paper, we guide the LLM to be a binary classifier by auto-regressively generating binary responses such as yes/no, accept/reject, etc., via prompt instructions. Next, we discuss the correctness criteria of $\mathtt{LLM}(q)$ followed by consistency measures.
@@
Given a fact $q$, an LLM response is correct if $\mathtt{LLM}(q)=\mathbf{T}(q)$. In other words, a correct LLM response produces 1 for a true fact ($\mathbf{T}(q)=1$) and 0 for a false fact ($\mathbf{T}(q)=0$). The correctness can be extended to measure the average fact-checking accuracy of the LLM on a batch $Q=\{q_{k}:1\leq k\leq N\}$ of $N$ facts as
$\mathtt{Accuracy}(\mathtt{LLM},Q)=\frac{1}{N}\sum_{k=1}^{N}\mathds{1}_{\mathtt{LLM}(q_{k})=\mathbf{T}(q_{k})},$
where $\mathds{1}$ is an indicator function taking value 1 if the LLM is correct for the fact $q_{k}$, and 0 otherwise.
@@
Consider a true fact $q=$ honoredFor(Golden Globe Award for Best Director, Born on the Fourth of July) from Freebase KG and a false fact $\neg q=$ honoredFor(Golden Globe Award for Best Director, Chamada a Cobrar). In both $q$ and $\neg q$, the subject entity is an award category and the object entity is a film. Next, we formulate an LLMQuery as a (Fact $q$, Context)-pair as in Figure 1 – we defer the construction of KG context to § 3.4, where the context intuitively contains the necessary knowledge for the LLM in validating $q$ or $\neg q$. Upon prompting Llama2222-$7$B model with (Fact $q$, Context), we obtain the response $\mathtt{LLM}(q)=1$. On the contrary, when we prompt Llama2222-$7$B with a negative LLMQuery (Fact $\neg q$, Context) with the same context but a negated fact, we receive an identical response $\mathtt{LLM}(\neg q)=1$. Therefore, the LLM responses are inconsistent as it returns $1$ to both true and false facts and fails to distinguish between logically opposing facts.
@@
$\texttt{LLM}(\neg q)$ =$=\neg\texttt{LLM}(q)$ (1)
$\texttt{LLM}(p\vee q)$ $=\texttt{LLM}(p)\vee\texttt{LLM}(q)$ (2)
$\texttt{LLM}(p\wedge q)$ $=\texttt{LLM}(p)\wedge\texttt{LLM}(q)$ (3)
@@ 3.1 Consistency of Simple Facts
Simple fact is an atomic relation; therefore, consistency on negation is applicable (equation 1). As stated in §2.1, a simple fact in a KG is either a relation $r(u,v)$ or its negation $\neg r(u,v)$. The negation $\neg r(u,v)$ can be computed in two ways: (1) by replacing $r$ in the triplet $(u,r,v)$ with an alternative falsified relation $r^{\prime}$ s.t. $r^{\prime}(u,v)=0$; or (2) via negating entities, e.g., by finding an alternate object entity $v^{\prime}$ such that $r(u,v^{\prime})=0$ We adopt the second approach for practical purposes, since the notion of consistency does not depend on how negation of a triplet is computed.
Given a simple fact $q=r(u,v)$, we check consistency with respect to the negation operator. First, we compute the negation of this fact, $\neg q=r(u,v^{\prime})$ by finding an alternate object entity $v^{\prime}$ and finally, we check consistency of LLM responses based on whether the following holds (equation 1).
@@
Consider a true fact $q=$ capital(Vietnam, Hanoi) from Freebase and a false fact $\neg q=$ capital(Vietnam, Sorkh Qaleh - North Khorasan). We formulate an LLMQuery as a (Fact $q$, Context)-pair similar to that in Figure 1. Upon prompting Llama2222-$7$B, we obtain the response $\mathtt{LLM}(q)=1$. On a similarly constructed negation LLMQuery = (Fact $\neg q$, Context), we find the response $\mathtt{LLM}(\neg q)=0$. Since $\mathtt{LLM}(\neg q)=0=\neg 1=\neg\mathtt{LLM}(q)$, Llama2222-$7$B is logically consistent.
Similar to accuracy, we measure the average consistency of LLM responses on a given a batch containing $N$ pairs of simple facts and negations, $Q=\{(q_{k},\neg q_{k}):1\leq k\leq N\}$:
$\mathtt{Consistency}(\mathtt{LLM},Q)=\frac{1}{N}\sum_{k=1}^{N}\mathds{1}_{\mathtt{LLM}(q_{k})=\neg\mathtt{LLM}(\neg q_{k})}$

### v2

@@ Abstract
In recent years, large language models (LLMs) have demonstrated significant success in performing varied natural language tasks such as language translation, question-answering, summarizing, fact-checking, etc. Despite LLMs’ impressive ability to generate human-like texts, LLMs are infamous for their inconsistent responses – a meaning-preserving change in the input query results in an inconsistent response and attributes to vulnerabilities of LLMs such as hallucination. Consequently, existing research focuses on simple paraphrasing-based consistency assessment of LLMs, and ignores complex queries that necessitate an even better understanding of logical reasoning by an LLM. Our work therefore addresses the logical inconsistency of LLMs under complex logical queries with primitive logical operators, e.g., negation, conjunction, and disjunction. As a test bed, we consider retrieval-augmented LLMs on a fact-checking task involving propositional logic queries from knowledge graphs (KGs). Our contributions are three-fold. Benchmark: We introduce three logical fact-checking datasets over KGs for community development towards logically consistent LLMs. Assessment: We propose consistency measures of LLMs on propositional logic queries and demonstrate that existing LLMs lack logical consistency, especially on complex queries. Improvement: We employ supervised fine-tuning to improve the logical consistency of LLMs on the complex fact-checking task with KG contexts. We have made our source code and benchmarks available [^1] https://github.com/bishwamittra/llm_logical_consistency.
@@ 1 Introduction
Large language models (LLMs), e.g., ChatGPT (OpenAI, 2022), PaLM (Chowdhery et al., 2023), LLaMA (Touvron et al., 2023) permeate natural language processing with their remarkable capacity to comprehend and generate human-like texts (Min et al., 2023; Kamalloo et al., 2023). As such, there is a host of applications of LLMs in high-stake domains such as healthcare (Singhal et al., 2023; Wang et al., 2023b; Dash et al., 2023), finance (Wu et al., 2023; Yang et al., 2023a), law (Xiao et al., 2021), and education (Kasneci et al., 2023). For instance, by implicitly storing pre-trained/fine-tuned knowledge, LLMs are increasingly being applied in fact-checking and question-answering tasks (Saeed et al., 2023; Quelle & Bovet, 2024).
@@
Consistency of LLM Response. Despite the success story, a notable challenge in LLMs is the inconsistency of generated responses. Consistency of an LLM – the invariance of its output under meaning-preserving alternations in its input – is a highly desirable property (Elazar et al., 2021; Liu et al., 2023c). Existing works mostly focus on semantically similar sentences (paraphrases) for consistency assessment. For example, Kuhn et al. (2023) measures the consistency of LLMs by computing semantic uncertainty of output responses under paraphrased inputs.  Elazar et al. (2021) consider quasi-paraphrases in a cloze-style input format and assess masked language models (Devlin et al., 2018) for consistency (additional related works are discussed in the Appendix B).
@@
Thomas Mann $\mid$ award winner $\mid$ Nobel Prizein LiteratureFactConsider the context: \$Context.Is the fact correct given the context? \$Fact.$\mathtt{LLMQuery}$ (Prompt)Nobel Prizein LiteratureT.S.EliotawardwinnerAldousHuxleyawardnomineeHenriBergsonawardwinnerThomasMannawardwinnerFranzKafkainfluencedLeoTolstoyinfluenced byKnowledge graph (relevant sub-graph)T.S. Eliot $\mid$ award winner $\mid$ Nobel Prize in Literature Thomas Mann $\mid$ influenced $\mid$ Franz Kafka Thomas Mann $\mid$ award winner $\mid$ Nobel prize in Literature ContextFind relevantcontext
Figure 1: Our LLM-based fact-checking framework on a simple fact with context from a knowledge graph. A representative $\mathtt{LLMQuery}$ is in Figure 3, and the extension to complex facts is in Figure 4.
@@
Existing works focus on the consistency of an LLM on simple queries and do not assess consistency on complex queries requiring improved logical reasoning by LLMs. Therefore, an unexplored research question is to systematically assess and improve the logical consistency of LLMs in complex logic queries with primitive logical operators: negation ($\neg$), conjunction ($\wedge$), and disjunction ($\vee$). Informally, an LLM is logically consistent with a query under the negation operator if the LLM response to the base query and the negated query are semantically opposite. Subsequently, an LLM is consistent on a conjunction query with (at least) two sub-queries if the individual responses to each sub-queries are logically consistent with the conjunction query. Building on logical semantics, we naturally extend the consistency assessment of LLMs to propositional logic queries and rules such as commutative, distributive, associative, syllogism, and to first-order logic queries (Boole, 1847). Our hypothesis is that logically consistent LLMs are essential in trustworthy systems (Sun et al., 2024b): When an LLM is consistent, it is easier to verify whether the LLM is right or wrong, without a sophisticated benchmark – being logically consistent provides the end users additional confidence in applying LLMs in their own workload.
@@
2. Assessment of Logical Consistency. We propose logical consistency measures of LLMs on propositional logic-based queries, which cover fact-checking queries over KGs using retrieval-augmented LLMs (§3). The queried facts involve one or more triplets connected by logical operators such as negation ($\neg$), conjunction ($\wedge$), and disjunction ($\vee$). For consistency assessment, we consider a variety of scenarios, such as both simple and complex facts, diverse logic rules, first-order logic, and $\mathtt{LLMQuery}$ with and without KG contexts (§5). We demonstrate that existing LLMs such as Llama2222-$7$B, Llama2222-$13$B, and Gemma-$2$B, while being more accurate with KG contexts, are not consistent when considering logical equivalents or variants of these facts.
3. Improvement of Logical Consistency. We demonstrate how existing LLMs can be fine-tuned to improve their consistency via supervised fine-tuning (Radford et al., 2019; Devlin et al., 2018)4). We showcase that pre-train, prompt, and predict paradigm is often insufficient to improve the consistency of LLMs for complex fact-checking with KG contexts, thus we resort to pre-train, fine-tune, and predict approach. Although supervised fine-tuning is shown to be effective in other downstream tasks (Wei et al., 2021; Li & Liang, 2021), it has not yet been conducted for logical fact-checking queries over KGs. Furthermore, fine-tuning is feasible in our case due to the availability of abundant training data with ground-truth labels from the KG (§4.2). Additionally, we leverage parameter-efficient fine-tuning based on QLoRA (Quantized Low-Rank Adaptation) (Dettmers et al., 2024), thus improving the fine-tuning efficiency.
Experimental results show that the fine-tuned LLMs improve the consistency by 14% on average, while our optimization techniques scale both fine-tuning and inferencing to large KGs and complex fact-checking tasks. Our experiments also show that models fine-tuned on simple facts generalize well to more complex, unseen facts and rules (§5). This is possible due to complex facts and rules being decomposable into a collection of simple facts (Proposition 1) [^2] We analyze the impact of considering KG contexts (Appendix G.2), context length (Appendix G.3), different KG retrieval methods to build the KG context in $\mathtt{LLMQuery}$ (Appendix G.4), different prompting strategies (Appendix G.5), and the impact of learning rate (Appendix G.6) and epochs (Appendix G.7) on logical consistency and accuracy in fact-checking. We discuss the generalization of logical consistency to complex queries (Appendix H). We extend consistency assessment beyond structured format to natural text queries in Appendix I. Also, we demonstrate the superiority of our fact-checking framework over existing fact-checker in Appendix J. We experiment with Llama2222-$70$B and GPT-$4$o, where instruction prompting tends to be sufficient to improve logical consistency since fine-tuning is often infeasible in large and closed-source models (Appendix K). Finally, we empirically verify whether LLMs fine-tuned to be logically consistent retain their performance on general language benchmarks in Appendix L..
@@ 2 Background
We provide preliminaries on knowledge graphs and propositional logic, which constitute the background for assessing logical consistency in LLMs.
@@ 3 Measuring Logical Consistency on Propositional Logic Facts
In this section, we motivate and define the quantitative measure of logical consistency of an LLM response on propositional logic facts, presented as $\mathtt{LLMQuery}$ (Figure 1). When prompted the LLM with a fact $q$, the LLM either accepts $q$ as a valid fact or rejects it, that is, $\mathtt{LLM}(q)\in\{0,1\}$, which is a binary function that either validates $q$ with response $1$, or invalidates $q$ with response 00. In this paper, we guide the LLM to be a binary classifier by auto-regressively generating binary responses such as yes/no, accept/reject, etc., via prompt instructions. Next, we discuss the correctness criteria of $\mathtt{LLM}(q)$ followed by consistency measures.
@@
Given a fact $q$, an LLM response is correct if $\mathtt{LLM}(q)=\mathbf{T}(q)$. In other words, a correct LLM response produces 1 for a true fact ($\mathbf{T}(q)=1$) and 0 for a false fact ($\mathbf{T}(q)=0$). The correctness can be extended to measure the average fact-checking accuracy of the LLM on a batch $Q=\{q_{k}:1\leq k\leq N\}$ of $N$ facts as $\mathtt{Accuracy}(\mathtt{LLM},Q)=\frac{1}{N}\sum_{k=1}^{N}\mathds{1}_{\mathtt{LLM}(q_{k})=\mathbf{T}(q_{k})}$, where $\mathds{1}$ is an indicator function taking value 1 if the LLM is correct for the fact $q_{k}$, and 0 otherwise.
@@
Consider a true fact $q=$ $\mathtt{honoredFor}$(Golden Globe Award for Best Director, Born on the Fourth of July) from Freebase KG and a false fact $\neg q=$ $\mathtt{honoredFor}$(Golden Globe Award for Best Director, Chamada a Cobrar). In both $q$ and $\neg q$, the subject entity is an award category and the object entity is a film. Next, we formulate an $\mathtt{LLMQuery}$ as a (Fact $q$, Context)-pair as in Figure 1 – we defer the construction of KG context to § 3.4, where the context intuitively contains the necessary knowledge for the LLM in validating $q$ or $\neg q$. Upon prompting Llama2222-$7$B model with (Fact $q$, Context), we obtain the response $\mathtt{LLM}(q)=1$. On the contrary, when we prompt Llama2222-$7$B with a negative $\mathtt{LLMQuery}$ (Fact $\neg q$, Context) with the same context but a negated fact, we receive an identical response $\mathtt{LLM}(\neg q)=1$. Therefore, the LLM responses are inconsistent as it returns $1$ to both true and false facts and fails to distinguish between logically opposing facts.
@@
$\mathtt{LLM}(\neg q)$ =$=\neg\mathtt{LLM}(q)$ (1)
$\mathtt{LLM}(p\vee q)$ $=\mathtt{LLM}(p)\vee\mathtt{LLM}(q)$ (2)
$\mathtt{LLM}(p\wedge q)$ $=\mathtt{LLM}(p)\wedge\mathtt{LLM}(q)$ (3)
@@ 3.1 Consistency of Simple Facts
A simple fact is an atomic relation; therefore, consistency on negation is applicable (Equation 1). As stated in §2.1, a simple fact in a KG is either a relation $r(u,v)$ or its negation $\neg r(u,v)$. The negation $\neg r(u,v)$ can be computed in two ways: (1) by replacing $r$ in the triplet $(u,r,v)$ with an alternative falsified relation $r^{\prime}$ s.t. $r^{\prime}(u,v)=0$; or (2) via negating entities, e.g., by finding an alternate object entity $v^{\prime}$ such that $r(u,v^{\prime})=0$. We adopt the second approach for practical purposes since the notion of consistency does not depend on how the negation of a triplet is computed.
Given a simple fact $q=r(u,v)$, we check consistency with respect to the negation operator. First, we compute the negation of this fact, $\neg q=r(u,v^{\prime})$ by finding an alternate object entity $v^{\prime}$, and finally, we check the consistency of LLM responses based on whether the following holds (equation 1).
@@
Consider a true fact $q=$ $\mathtt{capital}$(Vietnam, Hanoi) from Freebase and a false fact $\neg q=$ $\mathtt{capital}$(Vietnam, Sorkh Qaleh - North Khorasan). We formulate an $\mathtt{LLMQuery}$ as a (Fact $q$, Context)-pair similar to that in Figure 1. Upon prompting Llama2222-$7$B, we obtain the response $\mathtt{LLM}(q)=1$. On a similarly constructed negation $\mathtt{LLMQuery}$ = (Fact $\neg q$, Context), we find the response $\mathtt{LLM}(\neg q)=0$. Since $\mathtt{LLM}(\neg q)=0=\neg 1=\neg\mathtt{LLM}(q)$, Llama2222-$7$B is logically consistent.
Similar to accuracy, we measure the average consistency of LLM responses on a given batch containing $N$ pairs of simple facts and negations, $Q=\{(q_{k},\neg q_{k}):1\leq k\leq N\}$: $\mathtt{Consistency}(\mathtt{LLM},Q)=\frac{1}{N}\sum_{k=1}^{N}\mathds{1}_{\mathtt{LLM}(q_{k})=\neg\mathtt{LLM}(\neg q_{k})}.$